# Partner in Crime: Boosting Targeted Poisoning Attacks against Federated Learning

### v1

@@ I Introduction
Federated Learning (FL) [1, 2] has recently emerged as a scalable approach to generate Machine Learning (ML) models without the need to share privately owned data. Indeed, with FL, data can be kept in disjoint local sets, inherently offering better privacy for sensitive information. This feature makes FL a promising framework for a multitude of applications with dense device deployments at the edge, such as resource allocation for low-latency vehicle-to-vehicle (V2V) communications [3], mobile keyboard prediction by Google [4] and patient similarity analysis in healthcare [5, 6]. However, despite its performance scalability and intrinsic privacy benefits, FL remains susceptible to numerous privacy [7, 8, 9, 10, 11] and security attacks [12, 13, 14]. This is primarily due to the distributed nature of FL, which increases the exposed attack surfaces through malicious FL participants.
@@
The focus of this paper is on targeted poisoning attacks, an area that has been less explored compared to untargeted attacks in prior research. In practice, the strength and impact of existing poisoning attacks [12, 18, 14] are limited by the number of subverted FL clients, the presence of defenses [23, 24, 25, 26], and the attack capability. In data poisoning attacks, adversaries can only affect the model updates by contaminating the local data or inserting poisonous data. This limits the ability of adversaries to affect the training process. In model poisoning attacks, adversaries are allowed to modify the model weights and the training process. This is much more powerful when compared to data poisoning attacks. However, the divergence between malicious and benign updates can be readily detected by well-established defenses, such as the Byzantine-resilient methods Krum [23] and Median [26]. To evade these defenses, researchers introduced stealthy model poisoning attacks [12], which offer a trade-off between remaining undetected and improving attack success rates (ASRs). Unfortunately, being stealthy comes at a high cost, and in many cases, the ASRs of poisoning attacks can be significantly reduced [12]. Furthermore, the majority of untargeted poisoning attacks are not applicable in targeted scenarios due to the distinct objectives and attack strategies involved.
@@
In this paper, we propose a general pre-training approach to boost the ASRs of existing targeted poisoning attacks without affecting the stealthiness of original (vanilla) attacks. As a framework for Boosting Targeted Poisoning Attacks, BoTPA is able to breach the boundary between the source and target classes in the latent feature space. More specifically, BoTPA selects specific classes excluding the source and target classes, designated as intermediate classes, based on their model update contributions to the poisoned training process. An Amplifier set is then constructed using the existing samples from these intermediate classes. Subsequently, BoTPA intelligently crafts malicious soft labels to replace normal data labels in the Amplifier set. In contrast to one-hot encoding labels (hard labels), soft labels [27, 28] composed of fractional numbers have a higher entropy because there are more values that the label vector can take on. This increase in the density of values makes subtle modifications of the data labels much harder to detect. The total label modification mass can be more easily spread over the label components while ensuring a bound on individual component deltas.
Existing work [12, 29, 14] focuses exclusively on the direct relationship between source and target classes when launching a targeted poisoning attack. The main innovation of our work lies in indirectly altering the latent feature distributions of the source and target classes, as well as their decision boundary, by modifying the labels of intermediate classes. To the best of our knowledge, ours is the first work to consider intermediate classes as part of the attack procedure.
@@
Contributions: We present the design, implementation, and evaluation of BoTPA against FL models. BoTPA involves two critical steps: (1) selecting intermediate classes that generate model updates similar to those of the source class during the poisoned training process and subsequently forming the Amplifier set; (2) crafting malicious soft labels for the samples within the Amplifier set. In the initial stage of BoTPA, the attacker trains a surrogate model using datasets from malicious clients, where the source labels have been altered to the target labels. By collectively considering samples from each class, the attacker employs the Class Similarity from Contribution Degrees [30] metric to identify the intermediate classes. Subsequently, the Amplifier set is formed by incorporating data samples from these intermediate classes. Next, based on the Class Similarity from Latent Feature Distributions, the attacker designs soft labels for the Amplifier set to enhance the local poisoning impact. Finally, within the malicious clients, the data labels in the Amplifier set are modified to the crafted soft labels.
@@
We apply BoTPA to data poisoning attacks [14], model poisoning attacks by explicitly boosting updates [12], and stealthy model poisoning attacks [12]. Throughout the rest of this paper, unless specified otherwise, the term “poisoning attack” refers to targeted poisoning attacks. Our experiments are conducted across three datasets: FMNIST, CIFAR-10, and CH-MNIST, with a range of malicious client percentages from 5% to 30%. The experimental results demonstrate that BoTPA can significantly enhance the effectiveness of targeted poisoning attacks while maintaining stealthiness under the Byzantine-resilient defenses Krum [23] and Median [26], as well as the state-of-the-art defense Flame [31]. The summary of our main contributions is given below:
@@ II Background and Related Work
In this section, we present the fundamentals of the FL system, followed by an exploration of how it can be compromised by targeted and untargeted poisoning attacks, as well as the defenses against them.
II-A Model of FL System
We consider an FL system comprising a global server and $K$ local clients, denoted by $\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{K}\}$. The FL system under poisoning attacks is illustrated in Figure 1. In this system, the local clients collaborate to develop a well-trained model while keeping their data private from both the server and other participants. The FL model provides a function $f_{\boldsymbol{w}}:\boldsymbol{x}\rightarrow y$ to map inputs to labels. Clients perform all training operations locally, while the server is responsible for aggregating the local model updates and broadcasting the global model to the clients. The workflow of the FL system is outlined as follows.
  • Step 1: A global model $\boldsymbol{w}_{0}$ is initialized in the server and then broadcast to clients.
  • Step 2: Each local client calculates the model update $\boldsymbol{\delta}_{t}^{(k)}$ and sends it back to the server.
  • Step 3: The server applies the aggregation rule to the received updates. With FedAvg [32], the global update is $\boldsymbol{\delta}_{t}=\sum_{k=1}^{K}\frac{n^{(k)}}{\sum_{k=1}^{K}n^{(k)}}\boldsymbol{\delta}_{t}^{(k)}$, where $n^{(k)}$ is the size of the dataset in the $k$-th client. The updated global model is then computed as $\boldsymbol{w}_{t}=\boldsymbol{w}_{t-1}+\boldsymbol{\delta}_{t}$.
  • Step 4: The server broadcasts the updated global model to clients. Steps 2 to 4 are iteratively performed until convergence.
@@ II-A Model of FL System
II-B Untargeted Poisoning Attacks in FL system
Tolpegin et al. [14] introduced label-flipping attacks, which modify only the training data labels. Bhagoji et al. [12] proposed model poisoning attacks that employ an alternative minimization approach to maintain stealthiness. Fang et al. [13] formulated the attack as an optimization problem to craft malicious updates in the opposite direction of the normal global model update without an attack. Zhang et al. [33] presented PoisonGAN, a method based on generative adversarial networks to generate fake data samples resembling the data distribution of malicious clients. A more stealthy attack approach is backdoor attacks[15, 34, 16], where misclassifications are induced through triggers attached to data features. However, our boosting method focuses solely on label manipulation without feature modification. Thus, backdoor attacks with triggers are beyond the scope of this paper.
@@ II-B Untargeted Poisoning Attacks in FL system
II-C Targeted Poisoning Attacks in FL system
Based on the capabilities of different attackers, we detail the attack strategies for targeted data poisoning [14] and targeted model poisoning [12] attacks below.
@@ II-C1 Targeted data poisoning attack
Before training, the attacker maliciously modifies data labels from the source class to the target class [14]. The attacker is not allowed to directly alter the model weights or the loss function.
@@ II-C2 Targeted model poisoning attacks
Two types of model poisoning attacks are introduced below, serving as baselines in this paper and referred to as vanilla attacks. It is important to note that the initial step of both targeted attacks involves changing the source labels to the target labels. (1) Model poisoning attack by explicitly boosting updates [12]: Instead of sending the normal updates $\delta^{(k)}$ back, malicious clients apply a boosting factor $\lambda$ to them. This approach aims to mitigate the downscaling effect by the averaging aggregation, FedAvg, by reporting $\lambda\delta^{(k)}$. This attack is referred to as the explicit model poisoning attack for simplicity in this paper. (2) Stealthy model poisoning attack by alternating minimization [12]: This strategy aims to improve the attack’s stealthiness by considering two detection metrics—accuracy on validation data and weight update statistics. Accordingly, two specific loss terms are incorporated into the poisoning objective function: training loss over the benign dataset and the distance between malicious and benign updates.
@@ II-D Defenses
The Krum and Multi-Krum [23] defenses select either one or multiple updates with the smallest distances to other updates as benign updates for the next communication round. The Median [26] defense computes the coordinate-wise median of local updates as the new update. The Trimmed Mean [26] defense averages the remaining part of updates after removing the largest and smallest fractions. Bulyan[25] combines Krum with Trimmed Mean. Contra [37] identifies suspicious models that display higher similarities with other local models, and then adaptively reduces their learning rates and likelihood of being selected for the next training round. FLtrust[24] first computes the cosine similarity between the global update and local updates as the trust scores, and then uses the average of normalized local updates weighted by trust scores as the global update. Flare[38] designs the trust score based on the distance of penultimate layer parameters. FLDetector [39] designs a malicious client detection method based on update consistency, which is implemented prior to secure aggregations.
In addition, recent studies have demonstrated that differential privacy (DP), achieved by clipping and adding noise to local updates, can effectively enhance the model’s robustness against poisoning attacks [40, 34, 41]. Flame [31] constrains the amount of injected DP noise by employing three defense phases: filtering deviated model updates, clipping scaled-up updates, and finally adding suitable noise to the updates.
@@ III Threat Model
Adversarial Goal. The adversary’s goal is to manipulate the FL model to misclassify samples from the source class as the target class.
Adversarial Capabilities. We assume that only the modification of existing data labels on malicious clients is required, without injecting additional malicious data samples or making extra modifications to model weights. BoTPA can be applied to both data poisoning and model poisoning attacks without disturbing their original workflow. Additionally, backdoor attacks involving the attachment of triggers to images are not within the scope of this study, as we do not modify data features.
@@ IV-A System Design Overview
During both data poisoning and model poisoning attacks, the attacker compromises a small subset of clients in the FL system to manipulate the training dataset or the model. Malicious information flows from these local clients to the server during the training phase, ultimately causing the model to make incorrect predictions. Although the capabilities of attackers differ between data poisoning and model poisoning attacks, both types of attacks begin by modifying data labels. In these vanilla attacks, only the data labels of the source class are modified, without considering the indirect impact of other classes regarding targeted misclassification. To fill this gap, we propose the boosting method, BoTPA, which focuses on label manipulation in the pre-training stage. The BoTPA overview is illustrated in Figure 2, with the steps outlined below:
@@ IV-B Construction of Amplifier Set
Input Similarity from Contribution Degrees. For neural networks, two visually similar training images may have a very distinct influence on the training process, including the directions and magnitudes of model updates. We refer to the impact of input samples on the model updates as their contribution degrees to the training process. In order to evaluate these contribution degrees, the input similarity, $IS_{\text{contrib}}$, is defined based on the output difference upon a model update [30]. Considering two inputs, $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$, and the model weights $\boldsymbol{w}$. The metric $IS_{\text{contrib}}$ is derived through the following steps: (1) slightly modifying the output value for $\boldsymbol{x}$ by $\varepsilon$ through the addition of a minor update $\delta\boldsymbol{w}$ to the model weights; (2) observing the extent to which the output for $\boldsymbol{x}^{\prime}$ changes due to $\delta\boldsymbol{w}$; (3) using the difference in output changes as $IS_{\text{contrib}}$. In step (2), the change in the output for $\boldsymbol{x}^{\prime}$ is given by $\varepsilon\frac{\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}\left(\boldsymbol{x}^{\prime}\right)\cdot\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}(\boldsymbol{x})}{\left\|\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}(\boldsymbol{x})\right\|^{2}}$, where the kernel $\frac{\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}\left(\boldsymbol{x}^{\prime}\right)\cdot\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}(\boldsymbol{x})}{\left\|\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}(\boldsymbol{x})\right\|^{2}}$ represents the influence of $\boldsymbol{x}$ over $\boldsymbol{x}^{\prime}$ [^1] For more details on the derivation, please refer to [30]. Thus, the symmetric normalized input similarity between $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$ from contribution degrees is defined as
@@
It is worth noting that we construct the Amplifier set using intermediate classes rather than individual data samples due to the inherent uncertainty in machine learning training processes [42]. Utilizing intermediate classes helps stabilize the similarity scores between individual images from different classes.

### v2

@@ I Introduction
Federated Learning (FL) [1, 2] has recently emerged as a scalable approach to generate Machine Learning (ML) models without the need to share privately owned data. Indeed, with FL, data can be kept in disjoint local sets, inherently offering better privacy for sensitive information. This feature makes FL a promising framework for a multitude of applications with dense device deployments at the edge, such as resource allocation for low-latency vehicle-to-vehicle (V2V) communications [3], mobile keyboard prediction by Google [4], intrusion detection in IoT networks [5, 6] and patient similarity analysis in healthcare [7, 8]. However, despite its performance scalability and intrinsic privacy benefits, FL remains susceptible to numerous privacy [9, 10, 11, 12, 13, 14] and security attacks [15, 16, 17]. This is primarily due to the distributed nature of FL, which increases the exposed attack surfaces through malicious FL participants.
@@
The focus of this paper is on targeted poisoning attacks, an area that has been less explored compared to untargeted attacks in prior research. In practice, the strength and impact of existing poisoning attacks [15, 21, 17] are limited by the number of subverted FL clients, the presence of defenses [26, 27, 28, 29], and the attack capability. In data poisoning attacks, adversaries can only affect the model updates by contaminating the local data or inserting poisonous data. This limits the ability of adversaries to affect the training process. In model poisoning attacks, adversaries are allowed to modify the model weights and the training process. This is much more powerful when compared to data poisoning attacks. However, the divergence between malicious and benign updates can be readily detected by well-established defenses [26, 29]. To evade these defenses, researchers introduced stealthy model poisoning attacks [15], which offer a trade-off between remaining undetected and improving attack success rates (ASRs). Unfortunately, being stealthy comes at a high cost, and in many cases, the ASRs of poisoning attacks can be significantly reduced [15]. Furthermore, the majority of untargeted poisoning attacks are not applicable in targeted scenarios due to the distinct objectives and attack strategies involved.
Prior research [15, 30, 17] has focused exclusively on the direct relationship between the source and target classes when executing targeted poisoning attacks. The most intuitive strategy [17] in these scenarios involves mislabeling data samples from the source class as the target class. However, there are multiple factors affecting the decision boundary, with data distribution playing a significant role [31, 32]. Notably, an important observation is that poorly separated classes in the feature space result in less distinct decision boundaries. Exploiting this insight, an attacker can deliberately manipulate the feature space to increase overlap by modifying the source class and other classes. This raises two critical research questions: (1) how to identify the additional classes that significantly affect the decision boundary between the source and target classes, and (2) how to mislabel these classes to enhance the poisoning effect while minimizing the likelihood of detection.
@@
In this paper, we propose a general pre-training approach to boost the ASRs of existing targeted poisoning attacks without affecting the stealthiness of original (vanilla) attacks. As a framework for Boosting Targeted Poisoning Attacks, BoTPA is able to breach the boundary between the source and target classes in the latent feature space. More specifically, BoTPA selects specific classes excluding the source and target classes, designated as intermediate classes, based on their model update contributions to the poisoned training process. An Amplifier set is then constructed using the existing samples from these intermediate classes. This design allows BoTPA to enhance the attack’s effectiveness without relying solely on direct source-target modifications. Subsequently, BoTPA intelligently crafts malicious soft labels to replace normal data labels in the Amplifier set. In contrast to one-hot encoding labels (hard labels), soft labels [33, 34] composed of fractional numbers have a higher entropy because there are more values that the label vector can take on. This increase in the density of values makes subtle modifications of the data labels much harder to detect. The total label modification mass can be more easily spread over the label components while ensuring a bound on the individual component deltas. To the best of our knowledge, ours is the first work to leverage intermediate classes for launching a poisoning attack.
@@
Contributions: We present the design, implementation and evaluation of BoTPA against FL models. BoTPA involves two critical steps: (1) selecting intermediate classes that generate model updates similar to those of the source class during the poisoned training process and subsequently forming the Amplifier set; (2) crafting malicious soft labels for the samples within the Amplifier set. In the initial stage of BoTPA, the attacker trains a surrogate model using datasets from malicious clients. By collectively considering samples from each class, the attacker employs the Class Similarity from Contribution Degrees [35] metric to identify the intermediate classes. Subsequently, the Amplifier set is formed by incorporating data samples from these intermediate classes. Next, based on the Class Similarity from Latent Feature Distributions, the attacker designs soft labels for the Amplifier set to enhance the local poisoning impact. Finally, within the malicious clients, the data labels in the Amplifier set are modified to the crafted soft labels.
@@
We apply BoTPA to data poisoning attacks [17] and stealthy model poisoning attacks [15]. Throughout the rest of this paper, unless specified otherwise, the term “poisoning attack” refers to targeted poisoning attacks. Our experiments are conducted across three datasets: FMNIST, CIFAR-10, and CH-MNIST, with a range of malicious client percentages from 5% to 30%. The experimental results demonstrate that BoTPA can significantly enhance the effectiveness of targeted poisoning attacks while maintaining stealthiness under the Byzantine-resilient defenses Krum [26] and Median [29], as well as the state-of-the-art defense Flame [36]. The summary of our main contributions is given below:
@@ II Background and Related Work
In this section, we present the fundamentals of the FL system, followed by an exploration of how it can be compromised by poisoning attacks, as well as the defenses against them.
II-A FL System
We consider an FL system comprising a global server and $K$ local clients, denoted by $\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{K}\}$. In this system, the local clients collaborate to develop a well-trained model while keeping their data private from both the server and other participants. All training operations are performed locally by the clients, while the server is responsible for aggregating the local model updates. The workflow of the FL system is outlined as follows: (1) The server initializes a global model $\boldsymbol{w}_{0}$ and broadcasts to clients. (2) Each local client computes the model update $\boldsymbol{\delta}_{t}^{(k)}$ and sends it back to the server. (3) The server applies the aggregation rule to the received updates. With FedAvg [37], the global update is $\boldsymbol{\delta}_{t}=\sum_{k=1}^{K}\frac{n^{(k)}}{\sum_{k=1}^{K}n^{(k)}}\boldsymbol{\delta}_{t}^{(k)}$, where $n^{(k)}$ is the size of the dataset in the $k$-th client. The updated global model is then computed as $\boldsymbol{w}_{t}=\boldsymbol{w}_{t-1}+\boldsymbol{\delta}_{t}$. (4) The server broadcasts the updated global model to the clients. Steps (2) to (4) are repeated iteratively until the model converges.
@@ II-A FL System
II-B Untargeted Poisoning Attacks in FL systems
Tolpegin et al. [17] introduced label-flipping attacks, which modify only training data labels. Bhagoji et al. [15] proposed model poisoning attacks that employ an alternative minimization approach to maintain stealthiness. Fang et al. [16] formulated the attack as an optimization problem to craft malicious updates in the opposite direction of the normal global model update without an attack. Zhang et al. [38] presented PoisonGAN, a method based on generative adversarial networks to generate fake data samples resembling the data distribution of malicious clients. A more stealthy attack approach is backdoor attacks[39, 18, 40, 19], where misclassifications are induced through triggers attached to data features. However, our boosting method focuses solely on label manipulation without feature modification. Thus, backdoor attacks with triggers are beyond the scope of this paper.
@@ II-B Untargeted Poisoning Attacks in FL systems
II-C Targeted Poisoning Attacks in FL systems
Here, we outline the strategies for both targeted data poisoning [17] and targeted model poisoning [15] attacks, which serve as baseline adversarial methods in this study and are collectively referred to as vanilla attacks. Figure 1 provides an overview of the FL system under these poisoning attacks. Specifically, attacker 1 performs model poisoning and attacker 2 conducts data poisoning. Both aim to degrade the classification performance of the target model. While both attackers can manipulate the training data, the model poisoning attacker can further alter the FL model directly, leading to more severe performance degradation. A detailed discussion of these attack strategies is given below.
@@ II-C1 Targeted data poisoning attack
Before training, the attacker maliciously modifies the data labels from the source class to the target class [17]. The attacker is not allowed to directly alter the model weights or the loss function.
@@ II-C2 Targeted model poisoning attacks
Two types of targeted model poisoning attacks are introduced below. It is important to note that both attacks begin by altering the source labels to match the target labels. (1) Model poisoning attack by explicitly boosting updates [15]: Instead of sending the normal updates $\delta^{(k)}$ back, malicious clients apply a boosting factor $\lambda$ to them. This approach aims to counteract the downscaling effect caused by the FedAvg aggregation method by reporting $\lambda\delta^{(k)}$. (2) Stealthy model poisoning attack by alternating minimization [15]: This strategy aims to improve the attack’s stealthiness by considering two detection metrics—accuracy on validation data and weight update statistics. Accordingly, two specific loss terms are incorporated into the poisoning objective function: training loss over the benign dataset and the distance between malicious and benign updates.
@@ II-D Defenses
The Krum and Multi-Krum [26] defenses select either one or multiple updates with the smallest distances to other updates as benign updates for the next communication round. The Median [29] defense computes the coordinate-wise median of local updates as the new update. The Trimmed Mean [29] defense averages the remaining part of updates after removing the largest and smallest fractions. Bulyan[28] combines Krum with Trimmed Mean. The aforementioned defenses are Byzantine-resilient and straightforward to implement, as they are either non-parametric or require minimal parameter tuning. Advanced poisoning attacks, such as those presented in [15], have demonstrated the ability to compromise these defenses by manipulating model weights. However, to circumvent Byzantine-resilient mechanisms, these attacks require a trade-off that significantly reduces their success rates. Contra [43] identifies suspicious models that display higher similarities with other local models, and then adaptively reduces their learning rates and likelihood of being selected for the next training round. FLtrust[27] first computes the cosine similarity between the global update and local updates as the trust scores, and then uses the average of normalized local updates weighted by trust scores as the global update. Flare[44] designs the trust score based on the distance of penultimate layer parameters. FLDetector [45] designs a malicious client detection method based on update consistency, which is implemented prior to secure aggregations. Contra, FLTrust, FLARE, and FLDetector are effective in defense across various attack scenarios. However, these methods rely on different assumptions about the FL system and often require complex parameter tuning. For example, Contra heavily depends on detecting alignment (cosine similarity) among the updates from malicious clients. If the malicious local models exhibit diverse patterns, the effectiveness of this approach may be significantly reduced.
In addition, recent studies have demonstrated that differential privacy (DP) [46, 40, 47], implemented through clipping and adding noise to local model updates, can effectively enhance the FL model’s robustness against poisoning attacks. Among these studies, Flame [36] stands out as a particularly noteworthy approach. It constrains the amount of injected DP noise by employing three defense phases: filtering deviated model updates, clipping scaled-up updates, and finally adding suitable noise to the updates. In this paper, we include Flame as one of the baseline defense methods, as it has proven effective in mitigating the impact of poisoning attacks while preserving the model’s performance on the main task.
II-E Design and Application of Soft Labels
Traditionally, one-hot encoding, or hard label, has been widely used for training neural networks in classification tasks. However, hard labels fail to capture the shared underlying features across different classes. To address this problem, existing work has explored soft labels [48, 34, 49, 50], with a focus on enhancing the generalization and transferability of ML models by smoothing their decision boundaries. Two prominent methods are commonly used to generate soft labels. The first is label smoothing[34], which mixes hard labels with a uniform distribution over all classes. The second approach is knowledge distillation [51], where the knowledge from a larger and pre-trained model (teacher model) is transferred to a smaller model (student model). During this process, the output probabilities of the teacher model serve as soft labels for the student model, preserving the relative information between classes.
In this work, the attacker maliciously uses soft labels to mislabel intermediate classes, strategically influencing the decision boundary between the source and target classes in the latent feature space. This approach facilitates misclassification while preserving the stealthiness of the poisoning attacks.
@@ III Threat Model
Adversarial Goal. The adversary’s goal is to manipulate the FL model to cause samples from a specific source class, $c_{\text{src}}$, to be misclassified as a designated target class, $c_{\text{tgt}}$. Given a classification model $f_{\boldsymbol{w}}$, the attacker seeks to maximize the probability of incorrect classification:
$\max_{\boldsymbol{w}}\mathbb{P}\left(f_{\boldsymbol{w}}(\boldsymbol{x})=c_{\text{tgt}}\mid y=c_{\text{src}}\right),$ (1)
where $\boldsymbol{x}$ represents an input sample from the source class $c_{\text{src}}$ and $y$ denotes its true label.
Adversarial Capabilities. We assume that only the modification of existing data labels on malicious clients is required, without injecting additional malicious data samples or making extra modifications to model weights. BoTPA can be applied to both data poisoning and model poisoning attacks without disturbing their original workflow. Additionally, backdoor attacks involving the attachment of triggers to images are not within the scope of this study, as we do not modify data features. In terms of the percentage of compromised clients, we assume that the attacker can compromise fewer than 33% of the clients, adhering to the Byzantine fault tolerance model, which better represents real-world adversarial conditions. The more details on the target FL system and adversarial settings are given in Section V.
Defenses. We assume that established defenses against poisoning attacks, such as Byzantine-resilient aggregation techniques and DP-based mechanisms, are in place. However, this study does not consider adaptive defenses and we assume that the defender has no prior knowledge of the details of the proposed attack.
@@ IV-A System Design Overview
During both data poisoning and model poisoning attacks, the attacker compromises a small subset of clients in the FL system to manipulate the training dataset or the model. Although the capabilities of attackers differ between data poisoning and model poisoning attacks, both types of attacks begin by modifying data labels. In these vanilla attacks, only the data labels of the source class are modified, without considering the indirect impact of other classes regarding targeted misclassification. To fill this gap, we propose the boosting method, BoTPA, which focuses on label manipulation in the pre-training stage. The BoTPA overview is illustrated in Figure 2, with the steps outlined below:
@@ IV-B Construction of Amplifier Set
Input Similarity from Contribution Degrees. For neural networks, two visually similar training images may have a very distinct influence on the training process, including the directions and magnitudes of model updates. We refer to the impact of input samples on the model updates as their contribution degrees to the training process. In order to evaluate these contribution degrees, the input similarity, $IS_{\text{contrib}}$, is defined based on the output difference upon a model update [35]. Considering two inputs, $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$, and the model weights $\boldsymbol{w}$. The metric $IS_{\text{contrib}}$ is derived through the following steps: (1) slightly modifying the output value for $\boldsymbol{x}$ by $\varepsilon$ through the addition of a minor update $\delta\boldsymbol{w}$ to the model weights; (2) observing the extent to which the output for $\boldsymbol{x}^{\prime}$ changes due to $\delta\boldsymbol{w}$; (3) using the difference in output changes as $IS_{\text{contrib}}$. In step (2), the change in the output for $\boldsymbol{x}^{\prime}$ is given by $\varepsilon\frac{\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}\left(\boldsymbol{x}^{\prime}\right)\cdot\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}(\boldsymbol{x})}{\left\|\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}(\boldsymbol{x})\right\|^{2}}$, where the kernel $\frac{\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}\left(\boldsymbol{x}^{\prime}\right)\cdot\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}(\boldsymbol{x})}{\left\|\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}(\boldsymbol{x})\right\|^{2}}$ represents the influence of $\boldsymbol{x}$ over $\boldsymbol{x}^{\prime}$. [^1] For more details on the derivation, please refer to [35]. Thus, the symmetric normalized input similarity between $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$ from contribution degrees is defined as
@@
It is worth noting that we construct the Amplifier set using intermediate classes rather than individual data samples due to the inherent uncertainty in ML training processes [52]. Utilizing intermediate classes helps stabilize the similarity scores between individual images from different classes.